OpenStudy (anonymous):

16. y = 10 – 2x 2y = 5 – x either use addition or subtraction, can someone please show me how to solve this

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

y = 10 – 2x 2y = 5 – x -> y = 5/2 -x/2 ___________ 10-2x=2.5- 0.5x +2x +2x ___________ 10=2.5-1.5x -2.5-2.5 ___________ -7.5=1.5x /1.5 /1.5 ___________ 5=x ____________ y=10-2(5) y=0